6. Use Mathematical Induction to show that (21 - 1)(2i+1) n for all integers n > 1. 2n +1 (5 marks) i=1
Answer #1

In this question we have to use the mathematical induction so for his we have to use the following steps:

Step 1: In this prove the given statement for n=1,2,3.

Sep 2: In this suppose the statement is true for n=k for some integer k \geq 1

Then Step 3: By using the supposition in the Step 2 prove the statement for the n=k+1

Then we will have that this statement is true for all n \geq 1

Mathematical Use Induction to show that coll Cat - Dla for all integrunn no is ky, made in cise supposition 31 (Qirolaiti) an2. n = 1 for sell Lns I to I g Şumatim cis fuom so islo oney { Q-1 (i-ai+ in I (2-) LaH) 211)-1) (211)+1) ㅗ 3 Rus n Je L 211)for 3 3 m3 Lus (21-vjeariti) in + (2015-1)201341) * (ata)13 (212+7 ㅗ (2la)-1(2(a) +) (213)-1)(30)+) + L + 그 (6-1)(6+1) (2-1)(Proof! 5 RH mow consider M (ai-1) (ait) 1에 + -1) (2 H) merim 21R+1)-0(ak+1)+1 R by using @ ie M (2-1)(Qih) R ARH 키 R. + ㅗ (@k